The wrestling team leaves for Fargo traveling north on a slow bus at 57mph at the same time as the baseball team leaves for Texas traveling south at 74mph. How long until they are 400 miles apart?
How do you work it out? What's the formula that's the thing I don't understand with story problems.

Answer by checkley71(8403)

SEEING AS THE BUSES ARE TRAVELING IN OPOSITE DIRECTION THEN THEIR TWO MILAGES ARE ADDITIVE.
57X+74X=400
131X=400 (THEIR COLLECTIVE MILEAGE IS 131 MPH).
X=400/131
X=3.053435 HOURS IS THE TIME AFTER THEY LEAVE THAT THEY WILL BE 400 MILES APART.
PROOF:
57*3.053435+74*3.053435=400
174.0458+225.95419=400
400=400